# VMem: Consistent Interactive Video Scene Generation with Surfel-Indexed View Memory

### v1

@@ 3.2 New View Generation
Specifically, we base our generator on the recently released pre-trained SEVA [46] model. Given the plug-and-play nature of our memory module, VMem can work with other image-set generators as well. While our model can operate with SEVA out-of-the-box, we find that fine-tuning it for our task further improves the efficiency and quality of the generated images. This is due to two key reasons.
First, SEVA is trained to work with a large fixed total number of reference and target frames, which requires high computational resources unavailable to most. Our memory-efficient design selects only the most relevant reference views from all possible context views, leading to better efficiency. In addition, the interactive nature of our task requires generating only a small number of new frames at a time, before the user inputs new commands. Thus, we fine-tune the model to work with a smaller fixed number of reference and target views.
Second, unlike SEVA’s original setting where the target views may be spatially sparse, our target views follow a coherent trajectory with smoothly varying camera poses. This structured setting simplifies the generation task and enables the model to benefit from domain-specific adaptation through fine-tuning. Additional training details are provided in Sec. 4.1.
@@
View Retrieving Approaches LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\downarrow$ $R_{\text{dist}}$downarrow $T_{\text{dist}}$downarrow
Temporal 0.194 7.52 0.018 1.942 0.458
Field of View (FOV) 0.424 13.11 0.192 1.782 0.285
VMem 0.371 14.82 0.275 0.793 0.124
Table 4: Ablation study. LPIPS (lower is better), PSNR (higher is better), SSIM (higher is better), $R_{\text{dist}}$ (lower is better), $T_{\text{dist}}$ (lower is better). The setting without Surfel-Indexed Memory uses the $K$ latest views to generate the novel view.
@@ 4.6 Ablation Study
We also ablate our memory module design on the cycle-trajectory setting to validate the effectiveness of the proposed components of VMem. Specifically, we compare the performance of VMem against temporal-only and field-of-view-based (FOV) retrieval. For temporal-only, we do not use any memory mechanism and simply condition the novel view generator with the most recent $K$ views. For the FOV-based memory mechanism, we select the top-$K$ context views that have the highest field-of-view overlap with the target view. Specifically, we project a uniform grid of 3D points from the target camera frustum and compute the number of points visible from each past view based on depth consistency and camera intrinsics. The FOV overlap score is defined as the proportion of target-view frustum points that are also visible from the candidate context view. We rank all past views by this score and retrieve the top-$K$ views with the highest overlap, thus encouraging the selection of views that share the most visual coverage with the target frame.
As shown in Tab. 4, Surfel-Indexed View Memory achieves the best performance. These results highlight the effectiveness of our memory design in capturing long-range dependencies and maintaining consistency across complex camera trajectories.

### v2

@@ 3.2 Novel View Generator
Specifically, we base our generator on the recently released pre-trained SEVA [46] model. Given the plug-and-play nature of our memory module, VMem can work with other image-set generators as well. While our model can operate with SEVA out-of-the-box, the original SEVA is trained with a large fixed total number of reference and target frames ($K+M=21$), which demands substantial computational resources that may be prohibitive for many applications. To address this, we fine-tune a more efficient version that operates with a reduced number of reference frames ($K=4$) and target views ($M=4$) and use it for our main experiments. Our ablation studies (Sec. 4.6) demonstrate that this fine-tuned version, when combined with VMem, achieves comparable performance to the original SEVA while delivering approximately 10$\times$ speedup at inference time.
@@
Context views $K$ View Retrieval LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ $R_{\text{dist}}$downarrow $T_{\text{dist}}$downarrow
17 Temporal 0.477 13.92 0.188 0.976 0.254
Camera Distance (SEVA [46]) 0.397 15.72 0.297 0.821 0.392
Field of View 0.374 15.75 0.292 0.911 0.382
VMem 0.304 18.15 0.377 0.892 0.165
4 Temporal 0.794 7.52 0.018 1.942 0.458
Camera Distance 0.422 13.27 0.187 1.787 0.319
Field of View 0.424 13.11 0.192 1.782 0.285
VMem 0.381 14.82 0.275 0.793 0.124
Table 4: Ablation study on cycle trajectories from RealEstate10K. We compare different view retrieval strategies for two context view counts $K$. $K=17$ is the maximum context size in SEVA [46] for generating $M=4$ target views, while we use $K=4$ in our main experiments due to the $\sim$12$\times$ faster inference. The original SEVA strategy relies on camera distance for retrieval.
@@ 4.6 Ablation Study
We conduct an ablation study to assess the effectiveness of our proposed surfel-based memory indexing for context view retrieval in VMem. Specifically, we evaluate our approach in the cycle-trajectory setting and compare it against alternative view retrieval strategies: temporal-based, camera distance-based, and field of view-based retrieval. For temporal-based retrieval, we select the most recent $K$ views from memory. For camera distance-based retrieval, we select the top $K$ views with cameras closest to the target view camera—this strategy is also employed by SEVA [46], which serves as our backbone novel view generator. For field of view-based retrieval, we select the top $K$ views with the highest field of view overlap with the target view.
We also evaluate the performance trade-off between our lightweight, fine-tuned novel view generator using $K=4$ context views versus the original SEVA configuration with $K=17$ context views. The results are reported in Tab. 4.
Our findings demonstrate that VMem consistently improves the performance of the novel view generator across both context view configurations. The improvement is most pronounced when using fewer context views ($K=4$), highlighting the effectiveness of our surfel-based memory indexing.
Notably, our VMem with only $K=4$ views recovers most of the performance achieved with $K=17$ views in the original SEVA setting while delivering approximately 12×12$\times$12 × speed improvement. On an RTX 4090 GPU, our lightweight novel view generator takes 4.2 seconds per frame, while the original SEVA takes 50 seconds per frame.